# Comparison of the Effects of Interaction with Intentional Agent and Artificial Intelligence using fNIRS

### v1

@@ Abstract
As societal interactions increasingly involve both intentional and unintentional agents, understanding their effects on human cognition becomes paramount. This study investigates the neural correlates of interacting with intentional versus artificial agents in a simulated tennis game scenario. Employing functional near-infrared imaging spectroscopy (fNIRS), we analyzed brain activity in 50 male participants during gameplay against both types of opponents. Our methodological approach ensures ecological validity by simulating real-world decision-making scenarios while participants undergo fNIRS scanning, avoiding the constraints of traditional neuroimaging methods. We focus on six prefrontal cortex channels, leveraging the 10-20 system, to capture nuanced differences in brain activity. Utilizing wavelet analysis, we dissected the data into frequency-specific differences, revealing subtle variations across different channels and frequency bands. Moreover, we quantified activity by comparing average data signals between rest and play modes across all points. Our findings unveil significant differences in neural activation patterns, particularly in one specific channel and frequency range, suggesting distinct cognitive processing when interacting with intentional agents. These results align with previous neuroimaging studies and contribute to understanding the neural underpinnings of human-agent interactions in naturalistic settings. While acknowledging study limitations, including sample homogeneity and spatial accuracy constraints, our findings underscore the potential of fNIRS in exploring complex cognitive phenomena beyond laboratory confines.
@@ Introduction
As our ability to mimic and reproduce human behaviors through technological advancements continues to advance, the frequency of our interactions with fellow human agents appears to be diminishing. The changing landscape of our lifestyle, greatly impacted by these technological advancements, suggests a future where artificial intelligence (AI) [1, 2, 3, 4] could marginalize certain aspects of human communication, pushing them to the outskirts of daily life [5, 6, 7, 8].
@@
One notable example of this shift is the impact of Large Language Models (LLM) [9, 10, 11, 12, 13] and computer vision [14, 15] technologies. These models are not only integrated into practical tools like voice-activated assistants but also take on roles resembling colleagues, consultants, or trainers. Consequently, our interactions extend beyond traditional human exchanges to include engagements with non-human entities, evolving into social interactions, collaborative efforts, and decision-making instances [8, 16, 17, 18].
@@
As AI integrates more profoundly into our lives, a pivotal question emerges: How will a brain, finely tuned by evolution for interactions with intentional agents of the same species, adapt to the rapid changes in our social dynamics? Reflecting on Julian Jaynes’ insight that ’we are always assuming a space behind our companion’s eyes into which we are talking, similar to the space we imagine inside our own heads where we are talking from’ [19], we contemplate the consequences of heightened interaction with entities perceived to lack a mind or intention. In essence, what will be the impact on our daily lives and our brain? Considering common-sense psychology [20, 21] and evolutionary theories [22, 23, 24], we anticipate that different cognitive attitudes will manifest through distinct brain signals.
@@
In the realm of distinguishing between intentional and unintentional agents, recent strides in the field of theory of mind[25] have witnessed an emergence of comprehensive studies. As a matter of fact, theory of mind is the ability to understand that others have thoughts, beliefs, and feelings that may differ from one’s own [26, 27]. These investigations explore the intricate relationships governing these distinct mental states, drawing inspiration from prominent theoretical frameworks such as Dennett’s ’intentional stance’ theory [28, 29, 30]. Accordingly, a groundbreaking study, ’Imaging the Intentional Stance in a Competitive Game,’ conducted by Gallagher et al., utilized a PET scanner to measure brain activity in participants engaged in a game of rock-paper-scissors. The experimental conditions varied, including interactions with intentional and unintentional agents, as well as a randomization system. Notably, the study revealed a significant difference in activity detected in the anterior paracingulate cortex (aPCC) [31]. A parallel experiment, employing a similar protocol with slight modifications and utilizing an fMRI machine, reported bilateral activation in the medial prefrontal cortex (MPFC), temporoparietal junction (TPJ), and the right thalamus when contrasting participants playing against a human versus playing a computer generating moves at random [32]. Building on these findings, additional studies have explored a diverse array of games, such as the trust game, a punishment game, the Ultimatum Game (UG), and the Prisoner’s Dilemma Game [33, 34, 35, 36]. The overarching aim of these endeavors is to deepen our understanding of interactions with non-human agents [36].
@@
To build upon the findings of previous research and explore further, we designed an ecologically valid experiment using fNIRS, a portable method that enables us to identify brain activity in a real-life setting, probing to a depth of 2-3 cm based on hemoglobin’s optical properties [37]. Our study differs significantly in key aspects: our brain scanning method, uses another kind of task, and provides an atmosphere more like real life than a lab condition.
@@
Our primary goal is to ascertain whether there are discernible signs of these mental states (presence of intentional agent and artificial agent) in the prefrontal cortex after these changes. In contrast to earlier studies with strict and highly precise laboratory tasks, our research employs a simulated computer-based tennis game. This dynamic scenario requires participants to continuously predict their opponent’s actions and make rapid, calculated decisions, mirroring the cognitive demands of real-life situations characterized by repetitive and continuous decision-making. Our methodological refinement also ensures that participants feel at ease and respond more naturally, similar to real-world scenarios where they are not confined within a scanner, as is the case with PET and fMRI methods where subjects must remain stationary.
@@
In our study, participants engaged in two distinct game scenarios: one involving artificial intelligence (non-embodied) and the other featuring a human agent (embodied). To capture nuanced differences in brain activity in response to these scenarios, we chose to employ 6 channels based on the 10-20 system. Anticipating variations in activity across these channels, we recognize that this difference may not be uniform across all channels.
@@
The distribution of channels, with two in the lateral part and four in the frontal part of the prefrontal cortex, aligns with previous research highlighting the prefrontal cortex’s central role in this type of cognitive operation. Our limitation in obtaining signals from the upper areas of the head is attributed to the darkness of participants’ hair and the noise it introduced to the NIR signal. Nevertheless, areas with closer connections to Brodmann 9 and 10 are expected to exhibit a more pronounced difference in fNIRS signal when comparing the activity in these two different scenarios.
@@
Additionally, considering the nature of this method and the output signal, which is based on measuring the highly variable concentration of hemoglobin with and without oxygen in the brain blood, we decomposed the signal into its components using the wavelet [38, 39] method. This allowed us to analyze the different frequency bands of the signal and understand the changes. Our hypothesis was that, in addition to a specific channel, this difference could be observed in a specific frequency band.
@@
Our decision to focus on this method takes into account the acknowledged limitations of previous research methodologies. We aim to contribute valuable insights to the ongoing discourse on the neural underpinnings of intentional and unintentional agents, particularly in more naturalistic conditions. Additionally, we aspire to contribute to the development of the fNIRS method, emphasizing its application in understanding cognition in operational situations rather than laboratory settings.
@@
Our study included 50 right-handed male participants, with an average age of 24.94 years (SD = 3.77, ranging from 18 to 30). We took precautions to ensure the integrity of the study by excluding individuals with a history of psychiatric, neurological, or medical conditions, as well as those who had taken psycho-pharmacological medications within the last two months. Additionally, all participants met specific criteria: they possessed a basic knowledge of tennis, had prior experience playing computer-based games, and had no prior experience in professional tennis simulator gaming of this nature. These criteria were put in place to ensure that the participants’ skill levels during the task were relatively consistent.
Ethical considerations played a pivotal role in our research. The study received approval from the Research Ethics Committees of the Faculty of Psychology and Education, granted under the code IR.UT.PSYEDU.REC.1401.074 (Approval ID). The committee conducted a thorough review of the research protocol, affirming our unwavering dedication to upholding the highest ethical standards throughout all aspects of the study. This rigorous ethical evaluation underscores our commitment to conducting research in strict adherence to established ethical guidelines.
@@
The fNIRS system emitted light at the standard near-infrared wavelengths of 740 nm and 860 nm, allowing for hemoglobin spectroscopy to measure the relative concentration of hemoglobin both with and without oxygen. The signals were recorded at a sampling rate of 10 Hz to capture real-time brain dynamics during interactions.
We collected fNIRS data using a 6-channel OxyMon system provided by Artinis Medical Systems. The placement of optodes followed the well-established international 10-20 system [40] used for positioning EEG electrodes. This configuration ensured comprehensive coverage of the prefrontal regions, with a specific emphasis on the central medial prefrontal areas near FPz, which are roughly situated within Brodmann area 9 and 10 [41]. These central medial prefrontal regions are recognized for their involvement in processes related to mind reading [42].
Additionally, we employed four channels to focus on the lateral areas at F7 and F8, which correspond to Brodmann areas 47-45 [41, 43]. These lateral channels are associated with functions related to altruistic behavior and inhibition [44, 45]. We selected these specific brain regions due to their relevance in mentalizing and social cognition, as supported by prior research [46, 47]. To convert the coordinates from the 10-20 system to the MNI standard, we utilized fOLD [48], a software with a user interface within the MATLAB environment. The resulting coordinates of the channels is depicted in Fig.1.
@@
Ao tennis 2 game as mind reading task
In this study, we employed a tennis simulator computer game, which proved to be well-suited to our research objectives for several reasons. This game offered several advantages, such as ease of learning for the participants, the ability to be played by both single and two players, and the adaptability of game difficulty levels through simple adjustments, ensuring the engagement of the participants.
@@
Furthermore, this game demonstrated excellent flexibility for implementing a block design, a key feature for our experimental setup. What truly sets this game apart, however, is its alignment with the nature of tennis itself. Tennis is a dynamic sport, demanding simultaneous planning for the next shot while strategizing to counter the opponent’s moves. This intricate process necessitates the continuous engagement of cognitive modules related to theory of mind and complex calculations within the brain. The game environment is shown in Fig.2.
@@
Data analysis played a pivotal role in unraveling the complex neural dynamics underlying our study. The fNIRS data underwent a rigorous and comprehensive preprocessing phase, which aimed not only to enhance data quality but also to eliminate potential artifacts that could affect the validity of our findings. This section provides a detailed account of the steps taken during data analysis, from preprocessing to statistical analysis. You see the process
It is crucial to emphasize that our findings rely on a two-dimensional approach facilitated by the wavelet method. In this framework, one dimension corresponds to the chosen channel, while the other dimension corresponds to a specific frequency range.Fig.5.
@@
Using oxygenated blood data
In analyzing our fNIRS data, it’s essential to recognize the application of the Beer-Lambert law [51, 52], which correlates the absorption of light in a medium with the concentration of absorbing substances such as hemoglobin. In our study, the device we used has already applied Beer-Lambert’s law to the raw data, providing concentrations of both oxygenated and deoxygenated hemoglobin.In our research, we concentrated solely on oxygenated hemoglobin (oxyHb) data, despite having access to both oxyHb and deoxygenated hemoglobin information [53, 54]. This decision was driven by the direct link between oxyHb and increased neural activity in the brain, aligning with our research goals. We aimed to pinpoint active brain regions in response to our experimental conditions [55].
Moreover, the choice to prioritize oxyHb was grounded in established research practices, as oxyHb is a reliable indicator of brain engagement across various cognitive tasks. We followed the conventions of many researchers in our field, emphasizing oxyHb as a robust marker of neural activity [56, 57, 58, 59].
@@
A Wavelet-Based Approach for decompose signal
To enhance measurement accuracy and introduce additional depth to the data analysis, we applied the Daubechies wavelet [60, 61] method (number 10) at level 10, resulting in the segmentation of data into 10 distinct components.6 This wavelet-based Approach offers several advantages in processing fNIRS data, particularly in capturing both temporal and frequency-domain information simultaneously [62, 63, 64]. And since in cognitive research, it is widely acknowledged that higher frequency bands primarily capture noise and heart rate fluctuations, while lower frequencies may also contain noise, the frequency range from 0.01 to 0.1 Hz is typically associated with cognitive processes [65, 54]. .
Participant psychophysiological states can affect heart rate, breathing rate, blood pressure, and autonomic regulatory activity at both intra- and extra-cerebral levels. Spontaneous hemodynamic oscillations are related to physiological vasomotor regulations and breathing-related fluctuations, characterized by signals at specific frequencies such as heart rate ($\sim$ 1 Hz), breathing rate ($\sim$ 0.3 Hz), Mayer waves ($\sim$ 0.1 Hz), and very low frequency oscillations (< 0.04 Hz) [65, 66]. In light of this understanding, we focused our analysis on four specific frequency bands, comprising components 6, 7, 8, and 9. The choice of these bands was guided by several factors, including their relevance to cognitive processes implicated in our task, such as attention, prediction, and mind-reading. Additionally, these bands correspond to the dynamics of hemodynamic responses and physiological fluctuations, allowing us to effectively capture the neural activity associated with our experimental manipulations. We utilized the multiresolution signal analyzer toolbox in MATLAB for this analysis [^1] https://www.mathworks.com/help/wavelet/ref/signalmultiresolutionanalyzer-app.html. Further details are provided in the accompanying Fig.

### v2

@@ Abstract
As societal interactions increasingly involve both intentional and unintentional agents, understanding their effects on human cognition becomes paramount. This study investigates the neural correlates of interacting with intentional versus artificial agents in a simulated tennis game scenario. Employing functional near-infrared spectroscopy (fNIRS), brain activity in 50 male participants during gameplay against both types of opponents was analyzed. The used methodological approach ensures ecological validity by simulating real-world decision-making scenarios while participants undergo fNIRS scanning, avoiding the constraints of traditional neuroimaging methods. Six prefrontal cortex channels are focused on, leveraging the 10-20 system, to capture nuanced differences in brain activity. Wavelet analysis was utilized to dissect the data into frequency-specific differences, revealing subtle variations across different channels and frequency bands. Moreover, activity was quantified by comparing average data signals between rest and play modes across all points using Generalized Linear Model (GLM). The findings unveil significant differences in neural activation patterns, particularly in one specific channel and frequency range, suggesting distinct cognitive processing when interacting with intentional agents. These results align with previous neuroimaging studies and contribute to understanding the neural underpinnings of human-agent interactions in naturalistic settings. While acknowledging study limitations, including sample homogeneity and spatial accuracy constraints, the study’s findings underscore the potential of fNIRS in exploring complex cognitive phenomena beyond laboratory confines.
@@ Introduction
As ability to mimic and reproduce human behaviors through technological advancements continues to advance, the frequency of interactions with fellow human agents appears to be diminishing. The changing landscape of lifestyle, greatly impacted by these technological advancements, suggests a future where artificial intelligence (AI) [1, 2, 3, 4] could marginalize certain aspects of human communication, pushing them to the outskirts of daily life [5, 6, 7, 8].
@@
One notable example of this shift is the impact of Large Language Models (LLM) [9, 10, 11, 12, 13] and computer vision [14, 15] technologies. These models are not only integrated into practical tools like voice-activated assistants but also take on roles resembling colleagues, consultants, or trainers. Consequently, interactions extend beyond traditional human exchanges to include engagements with non-human entities, evolving into social interactions, collaborative efforts, and decision-making instances [8, 16, 17, 18].
@@
As AI integrates more profoundly into lives, a pivotal question emerges: How will a brain, finely tuned by evolution for interactions with intentional agents of the same species, adapt to the rapid changes in social dynamics? Reflecting on Julian Jaynes’ insight that ’we are always assuming a space behind our companion’s eyes into which we are talking, similar to the space we imagine inside our own heads where we are talking from’ [19], the consequences of heightened interaction with entities perceived to lack a mind or intention are contemplated. In essence, what will be the impact on daily lives and the brain? Considering common-sense psychology [20, 21] and evolutionary theories [22, 23, 24, 25], it is anticipated that different cognitive attitudes will manifest through distinct brain signals.
@@
In the realm of distinguishing between intentional and unintentional agents, recent strides in the field of theory of mind[26] have witnessed an emergence of comprehensive studies. As a matter of fact, theory of mind is the ability to understand that others have thoughts, beliefs, and feelings that may differ from one’s own [27, 28]. These investigations explore the intricate relationships governing these distinct mental states, drawing inspiration from prominent theoretical frameworks such as Dennett’s ’intentional stance’ theory [29, 30, 31]. Accordingly, a groundbreaking study, ’Imaging the Intentional Stance in a Competitive Game,’ conducted by Gallagher et al., utilized a PET scanner to measure brain activity in participants engaged in a game of rock-paper-scissors. The experimental conditions varied, including interactions with intentional and unintentional agents, as well as a randomization system. Notably, the study revealed a significant difference in activity detected in the anterior paracingulate cortex (aPCC) [32]. A parallel experiment, employing a similar protocol with slight modifications and utilizing an fMRI machine, reported bilateral activation in the medial prefrontal cortex (MPFC), temporoparietal junction (TPJ), and the right thalamus when contrasting participants playing against a human versus playing a computer generating moves at random [33]. Building on these findings, additional studies have explored a diverse array of games, such as the trust game, a punishment game, the Ultimatum Game (UG), and the Prisoner’s Dilemma Game [34, 35, 36, 37]. The overarching aim of these endeavors is to deepen the understanding of interactions with non-human agents [37].
@@
To build upon the findings of previous research and explore further, an ecologically valid experiment was designed using fNIRS, a portable method that enables us to identify brain activity in a real-life setting, probing to a depth of 2-3 cm based on hemoglobin’s optical properties [38]. This study differs significantly in key aspects: the applied brain scanning method, the use of a different kind of task, and the provision of an atmosphere more akin to real life than a laboratory condition.
@@
The primary goal is to ascertain whether there are discernible signs of these mental states (presence of intentional agent and artificial agent) in the prefrontal cortex after these changes. In contrast to earlier studies with strict and highly precise laboratory tasks, this research employs a simulated computer-based tennis game. This dynamic scenario requires participants to continuously predict their opponent’s actions and make rapid, calculated decisions, mirroring the cognitive demands of real-life situations characterized by repetitive and continuous decision-making. The proposed methodological refinement also ensures that participants feel at ease and respond more naturally, similar to real-world scenarios where they are not confined within a scanner, as is the case with PET and fMRI methods where subjects must remain stationary.
@@
In this study, participants engaged in two distinct game scenarios: one involving artificial intelligence (non-embodied) and the other featuring a human agent (embodied). To capture nuanced differences in brain activity in response to these scenarios, a decision was made to employ 6 channels based on the 10-20 system. Variations in activity across these channels are anticipated, recognizing that this difference may not be uniform across all channels.
@@
The distribution of channels, with two in the lateral part and four in the frontal part of the prefrontal cortex, aligns with previous research highlighting the prefrontal cortex’s central role in this type of cognitive operation. This paper’s limitation in obtaining signals from the upper areas of the head is attributed to the darkness of participants’ hair and the noise it introduced to the NIR signal. Nevertheless, areas with closer connections to Brodmann 9 and 10 are expected to exhibit a more pronounced difference in fNIRS signal when comparing the activity in these two different scenarios.
@@
Additionally, considering the nature of this method and the output signal, which is based on measuring the highly variable concentration of hemoglobin with and without oxygen in the brain hemodynamic, the signal is decomposed into its components using the wavelet [39, 40] transform. This allowed us to analyze the different frequency bands of the signal and understand the changes. Our hypothesis was that, in addition to a specific channel, this difference could be observed in a specific frequency band.
@@
Our decision to focus on this method takes into account the acknowledged limitations of previous research methodologies. Valuable insights are aimed to be contributed to the ongoing discourse on the neural underpinnings of intentional and unintentional agents, particularly in more naturalistic conditions. Additionally, the aspiration is to contribute to the development of the fNIRS method, with emphasis placed on its application in understanding cognition in operational situations rather than laboratory settings.
@@
In this study, 50 right-handed male participants with an average age of 24.94 years (SD = 3.77, ranging from 18 to 30) are participated. Precautions were taken to ensure the integrity of the study by excluding individuals with a history of psychiatric, neurological, or medical conditions, as well as those who had taken psycho-pharmacological medications within the last two months. Additionally, all participants met specific criteria: they possessed a basic knowledge of tennis, had prior experience playing computer-based games, and had no prior experience in professional tennis simulator gaming of this nature. These criteria were put in place to ensure that the participants’ skill levels during the task were relatively consistent.
Ethical considerations played a pivotal role in this research. The study received approval from the Research Ethics Committees of the Faculty of Psychology and Education, granted under the code IR.UT.PSYEDU.REC.1401.074 (Approval ID). The committee conducted a thorough review of the research protocol, affirming unwavering dedication to upholding the highest ethical standards throughout all aspects of the study. This rigorous ethical evaluation underscores commitment to conducting research in strict adherence to established ethical guidelines.
@@
fNIRS data were collected using a 6-channel OxyMon system provided by Artinis Medical Systems. It emits light at the standard near-infrared wavelengths of 740 nm and 860 nm, allowing for hemoglobin spectroscopy to measure the relative concentration of hemoglobin both with and without oxygen. The signals were recorded at a sampling rate of 10 Hz to capture real-time brain dynamics during interactions.
The placement of optodes followed the well-established international 10-20 system [41] used for positioning EEG electrodes. This configuration ensured comprehensive coverage of the prefrontal regions, with a specific emphasis on the central medial prefrontal areas near FPz, which are roughly situated within Brodmann area 9 and 10 [42]. These central medial prefrontal regions are recognized for their involvement in processes related to mind reading [43].
Additionally, four channels were employed to focus on the lateral areas at F7 and F8, which correspond to Brodmann areas 47-45 [42, 44]. These lateral channels are associated with functions related to altruistic behavior and inhibition [45, 46]. These specific brain regions were selected due to their relevance in mentalizing and social cognition, as reported by prior research [47, 48]. To convert the coordinates from the 10-20 system to the MNI standard, fOLD [49], a software with a user interface within the MATLAB environment, is used. The resulting coordinates of the channels is depicted in Fig.1.
@@
Tennis 2 Game as a Mind Reading Task
In this study, a tennis simulator computer game was employed, which proved to be well-suited to this research objectives for several reasons. This game offered several advantages, such as ease of learning for the participants, the ability to be played by both single and two players, and the adaptability of game difficulty levels through simple adjustments, ensuring the engagement of the participants.
@@
This game demonstrated excellent flexibility for implementing a block design, a key feature for experimental setup. What truly sets this game apart, however, is its alignment with the nature of tennis itself. Tennis is a dynamic sport, demanding simultaneous planning for the next shot while strategizing to counter the opponent’s moves. This intricate process necessitates the continuous engagement of cognitive modules related to theory of mind and complex calculations within the brain. The game environment is shown in Fig.2.
@@
Data analysis played a pivotal role in unraveling the complex neural dynamics underlying this study. The fNIRS data underwent a rigorous and comprehensive preprocessing phase, which aimed not only to enhance data quality but also to eliminate potential artifacts that could affect the validity of the findings. This section provides a detailed account of the steps taken during data analysis, from preprocessing to statistical analysis.
It is crucial to emphasize that this study’s findings rely on a two-dimensional approach facilitated by the wavelet method. In this framework, one dimension corresponds to the chosen channel, while the other dimension corresponds to a specific frequency range.Fig.5.
@@
Utilizing OxyHb and DeoxyHb Data for Enhanced fNIRS Analysis
In analyzing our fNIRS data, it’s essential to recognize the application of the Beer-Lambert law [52, 53], which correlates the absorption of light in a medium with the concentration of absorbing substances such as hemoglobin. The Beer-Lambert law provides a fundamental basis for quantifying changes in hemoglobin concentration, crucial for interpreting neural activity measured by fNIRS. Our study utilized a device that has already applied the Beer-Lambert law to the raw data, yielding concentrations of both oxygenated (oxyHb) and deoxygenated (deoxyHb) hemoglobin.
While some previous analyses focused solely on oxyHb data [54, 55, 56, 57, 58, 59], we have expanded our investigation to include deoxyHb information as well. This decision aligns with standardized procedures highlighted in Yücel et al.’s paper on best practices for fNIRS publications [56]. Incorporating both oxyHb and deoxyHb data is crucial for providing a comprehensive understanding of neural activity, particularly in scenarios where physiological interferences related to the task may occur. For instance, previous research has indicated that individuals may alter their breathing patterns during tasks such as video gaming [60]. By examining both oxyHb and deoxyHb responses, we aim to capture the complete hemodynamic response and provide a more nuanced interpretation of neural activity in response to our experimental conditions [61, 59].
@@
A Wavelet-Based Approach to Decompose Signal
To enhance measurement accuracy and introduce additional depth to the data analysis, we applied the Daubechies wavelet [62, 63] at 10 decomposition levels to the signal. This segmentation resulted in the division of the data into 10 distinct components. The wavelet-based approach offers several advantages in processing fNIRS data, particularly in capturing both temporal and frequency-domain information simultaneously [64, 65, 66].
Previous fNIRS studies have shown that the frequency band of 0.01 to 0.1 Hz is typically associated with cognitive processes such as attention, prediction, and mind-reading [55, 67, 68, 69, 70, 71], while spontaneous hemodynamic oscillations are related to physiological vasomotor regulation and breathing-related fluctuations. These are characterized by signals at specific frequencies, including heart rate ($\sim$ 1 Hz), breathing rate ($\sim$ 0.3 Hz), Mayer waves ($\sim$ 0.1 Hz), and very low-frequency oscillations (<0.04 Hz) [67, 72]. Therefore, our interest lay in analyzing four specific frequency bands: levels 6 (0.0779-0.157 Hz), 7 (0.039-0.0874 Hz), 8 (0.0195-0.0392 Hz), and 9 (0.00978-0.0196 Hz), which fall within our target range.
While previous studies did not restrict the frequency ranges of interest to this extent and typically employed band-pass filters, cognitive results consistently emerged within this frequency range. The multiresolution signal analyzer toolbox in MATLAB facilitated this analysis, leveraging its capabilities for wavelet decomposition and frequency band analysis  [^1] https://www.mathworks.com/help/wavelet/ref/signalmultiresolutionanalyzer-app.html. Further details and results of the analysis are provided in Fig.6.